OpenStudy (anonymous):

How do I find initial velocty, acceleration, and final velocity? I am given t = 2.4 s and d = 5.6 m

OpenStudy (anonymous):

With 2 dimensional kinematics we can plug in the values for d and t. x = vt + 1/2at^2

OpenStudy (anonymous):

We know initial v is v= d/t which gives us 5.6/2.4

OpenStudy (anonymous):

May be wrong, I'd check your answers. It's been a while

rvc (rvc):

whats your actual question we know that velocity= displacement/ time u have the values for both you will get velocity... after that use this equation : s=ut+1/2 a t^2 s=d t=2.4 u= velocity value which u got above so u get a after that v=u+at u= that velocity value(intial) a= acceleration which u found out t=2.4 u will get V (final velocity)

OpenStudy (michele_laino):

please, what are the meanings of \(d=5.6\) meters and \(t=2.4\) seconds?